A 500 grams of copper at 300 0 C and 350 grams lead at 550 0 C are put in 1200grams of aluminium container filled with one litre of water at 20 0 C. Solve for the final temperature reading.


1
Expert's answer
2021-05-17T12:38:38-0400

"c_1m_1(300-t)+c_2m_2(550-t)=c_3m_3(t-20)+c_4m_4(t-20)"


"385\cdot 0.5\cdot(300-t)+130\cdot 0.35\cdot(550-t)=920\cdot 1.2\cdot(t-20)+4218\cdot 1\cdot(t-20)\to"


"t=34\ (°C)" . Answer